# These $\altaffiliationThese authors contributed equally to this work.$

### v1

@@ 2.3 Enhancement of Quantum Sensing
Finally, we investigated whether the observed enhancement of FND and hBN PL inside cavities improves their quantum sensing performance. We first focused on FNDs. We employed a custom built wide-field quantum sensing microscope [52] (Supplementary Fig. S6 to acquire ODMR spectra for FNDs inside the cavities investigated in the previous section as a function of cavity resonance $\lambda_{res}$. See Supplementary Information for full experimental details. These devices were fabricated on Si wafer substrates, which strongly absorb microwaves. Hence, microwaves were delivered via a loop antenna from the top-mirror side of the cavity, which made a direct comparison between FNDs inside and outside the cavity challenging due to top Ag mirror MW absorption only being present on the cavity side. However, it allowed us to investigate the ODMR contrast for all cavity resonances. We find the highest ODMR contrast of 4.4% for the cavity with $\lambda_{res}=650$ nm, which gradually drops to approximately 2% for resonance wavelengths of 500 nm and 800 nm (see Supplementary Fig. S7).
We were unable to acquire ODMR spectra for the hBN thin-film cavities. In general, the positive ODMR contrast of the spin pair emitters in the hBN NPs investigated here is below 1% even under optimized MW delivery conditions [14, 53]. Hence, one possible explanation is that the MW field inside our cavities was too low to efficiently drive the hBN spin transitions. For the hBN emitters investigated here, the MW-induced spin mixing that is the basis of ODMR occurs in a metastable state [53]. Hence, another possible explanation is that the PL decay observed for most hBN NPs inside the cavity became so fast, i.e. the excited state lifetime so short, that transitions to the metastable state were very inefficient resulting in a negligible population of the metastable state. Lastly, the decrease in hBN PL stability over time increased the measurement noise in ODMR spectra.
@@
Based on these results, we focused on the FND cavities, selected the $\lambda_{res}=650$ nm cavity for further analysis and fabricated devices on quartz substrates that allow MW delivery and imaging through the substrate (Fig. 4a) and hence a direct comparison between FNDs inside and outside the cavity. We fabricated one device with 100 nm FNDs and one with 20 nm FNDs (Adamas Nanotechnologies, USA) to also investigate if the much dimmer PL, mostly from the NV0 charge state of the 20 nm FNDs [54] can be enhanced for ODMR-based quantum sensing. Reflectance spectra for these devices is shown in Supplementary Fig. S8. Figs. 4b and c show ODMR spectra for 100 nm and 20 nm FNDs, respectively, inside and outside the cavity, with the characteristic NV- resonance at 2.87 GHz. The ODMR contrast increases by a factor of 2.2 (100 nm FNDs) and 2.8 (20 nm FNDs) while the full-width half-maximum (FWHM), indicated by the black arrows, remains constant within $\pm 1$ MHz. The PL brightness $\Phi$ increases 3-fold for 20 nm FNDs inside the cavity, while it decreases by about half for the 100 nm FNDs. In both cases, however, we observe spectrally narrower PL in the NV- spectral region between 650 and 700 nm (20 nm FNDs) and 700 and 800 nm (100 nm FNDs) for FNDs inside the cavity (see Supplementary Fig. S9). Spectral narrowing likely plays a dominant role in the enhancement of ODMR contrast observed in both samples’ cavity regions and suggests that the NV coupling to the cavity can selectively enhance NV- in our cavities.
Based on these measurements, we calculated the FND magnetic field sensitivity using [55]
@@ 2.4 Purcell Enhancement in Fabry-Pérot Micocavities
where $\lambda_{res}$ is the cavity resonance wavelength, $n$ is the refractive index of the cavity medium and $V_{eff}$ is the effective mode volume in the cavity. In this case, we use an idealized approach and assume perfect longitudinal confinement in the planar cavity, with no lateral confinement, where the assumption $V_{eff}\simeq 0.125(\frac{\lambda}{n})^{3}$ is appropriate [56].
@@ 2.6 Cavity Enhancement Within a Gaussian Projection of the Cavity Mode
To extract an projection of the cavity mode on emission from the cavity, cavity emission spectra were analyzed by fitting a sum of two Gaussian functions,
@@
The PL spectra of hBN nanoparticles are usually non-uniform due to the heterogeneous distribution of luminescent centers, and their surrounding environments. Emission typically originates from intrinsic point defects such as vacancies, donor–acceptor pairs, and impurity complexes, each associated with distinct ZPLs and phonon sidebands. Theoretical predictions [57, 58] and large-scale statistical studies [59] have shown that hBN emitters can be grouped into several discrete spectral families within the $\sim 1.6-2.2$ eV range, indicating the coexistence of multiple defect types even within nominally identical samples. Additionally, local variations in dielectric screening, strain, and surface termination can lead to shifts in emission energies and differences in spectral line shapes. These effects are further influenced by particle size, morphology, and interactions with substrates or the surrounding environment.
Non-uniformity is also affected by extrinsic contributions from organic residues, including aromatic polycyclic hydrocarbons introduced during processing and have been reported to act as additional luminescent centers in some hBN samples [60]. These emitters contribute their own PL features and complicate the spectral response of the nanoparticles. Temporal instabilities such as blinking and bleaching, which are common in nanoscale fluorophores due to charge trapping or non-radiative recombination, further add to spectral variability. These intrinsic and extrinsic effects account for the observed unpredictability in the PL spectra of hBN nanoparticles and highlight the challenges of achieving reproducible optical performance.

### v2

@@ 2.3 Enhancement of Quantum Sensing
Finally, we investigated whether the observed enhancement of FND and hBN PL inside cavities improves their quantum sensing performance. We first focused on FNDs. We employed a custom built wide-field quantum sensing microscope Scholten et al. [2021] (Supplementary Fig. S6 to acquire ODMR spectra for FNDs inside the cavities investigated in the previous section as a function of cavity resonance $\lambda_{res}$. See Supplementary Information for full experimental details. These devices were fabricated on Si wafer substrates, which strongly absorb microwaves. Hence, microwaves were delivered via a loop antenna from the top-mirror side of the cavity, which made a direct comparison between FNDs inside and outside the cavity challenging due to top Ag mirror MW absorption only being present on the cavity side. However, it allowed us to investigate the ODMR contrast for all cavity resonances. We find the highest ODMR contrast of 4.4% for the cavity with $\lambda_{res}=650$ nm, which gradually drops to approximately 2% for resonance wavelengths of 500 nm and 800 nm (see Supplementary Fig. S7).
We were unable to acquire ODMR spectra for the hBN thin-film cavities. In general, the positive ODMR contrast of the spin pair emitters in the hBN NPs investigated here is below 1% even under optimized MW delivery conditions Singh et al. [2025], Scholten et al. [2024]. Hence, one possible explanation is that the MW field inside our cavities was too low to efficiently drive the hBN spin transitions. For the hBN emitters investigated here, the MW-induced spin mixing that is the basis of ODMR occurs in a metastable state Scholten et al. [2024]. Hence, another possible explanation is that the PL decay observed for most hBN NPs inside the cavity became so fast, i.e. the excited state lifetime so short, that transitions to the metastable state were very inefficient resulting in a negligible population of the metastable state. Lastly, the decrease in hBN PL stability over time increased the measurement noise in ODMR spectra.
@@
Based on these results, we focused on the FND cavities, selected the $\lambda_{res}=650$ nm cavity for further analysis and fabricated devices on quartz substrates that allow MW delivery and imaging through the substrate (Fig. 4a) and hence a direct comparison between FNDs inside and outside the cavity. We fabricated one device with 100 nm FNDs and one with 20 nm FNDs (Adamas Nanotechnologies, USA) to also investigate if the much dimmer PL, mostly from the NV0 charge state of the 20 nm FNDs Wilson et al. [2019] can be enhanced for ODMR-based quantum sensing. Reflectance spectra for these devices is shown in Supplementary Fig. S8. Figs. 4b and c show ODMR spectra for 100 nm and 20 nm FNDs, respectively, inside and outside the cavity, with the characteristic NV- resonance at 2.87 GHz. The ODMR contrast increases by a factor of 2.2 (100 nm FNDs) and 2.8 (20 nm FNDs) while the full-width half-maximum (FWHM), indicated by the black arrows, remains constant within $\pm 1$ MHz. The PL brightness $\Phi$ increases 3-fold for 20 nm FNDs inside the cavity, while it decreases by about half for the 100 nm FNDs. In both cases, however, we observe spectrally narrower PL in the NV- spectral region between 650 and 700 nm (20 nm FNDs) and 700 and 800 nm (100 nm FNDs) for FNDs inside the cavity (see Supplementary Fig. S9). Spectral narrowing likely plays a dominant role in the enhancement of ODMR contrast observed in both samples’ cavity regions and suggests that the NV coupling to the cavity can selectively enhance NV- in our cavities.
Based on these measurements, we calculated the FND magnetic field sensitivity using Rondin et al. [2014]
@@ 2.4 Purcell Enhancement in Fabry-Pérot Micocavities
where $\lambda_{res}$ is the cavity resonance wavelength, $n$ is the refractive index of the cavity medium and $V_{eff}$ is the effective mode volume in the cavity. In this case, we use an idealized approach and assume perfect longitudinal confinement in the planar cavity, with no lateral confinement, where the assumption $V_{eff}\simeq 0.125(\frac{\lambda}{n})^{3}$ is appropriate Gérard [2003].
@@ 2.6 Cavity Enhancement Within a Gaussian Projection of the Cavity Mode
To extract a projection of the cavity mode on emission from the cavity, cavity emission spectra were analyzed by fitting a sum of two Gaussian functions,
@@
The PL spectra of hBN nanoparticles are usually non-uniform due to the heterogeneous distribution of luminescent centers, and their surrounding environments. Emission typically originates from intrinsic point defects such as vacancies, donor–acceptor pairs, and impurity complexes, each associated with distinct ZPLs and phonon sidebands. Theoretical predictions Sajid et al. [2018], Auburger and Gali [2021] and large-scale statistical studies Islam et al. [2024] have shown that hBN emitters can be grouped into several discrete spectral families within the $\sim 1.6-2.2$ eV range, indicating the coexistence of multiple defect types even within nominally identical samples. Additionally, local variations in dielectric screening, strain, and surface termination can lead to shifts in emission energies and differences in spectral line shapes. These effects are further influenced by particle size, morphology, and interactions with substrates or the surrounding environment.
Non-uniformity is also affected by extrinsic contributions from organic residues, including aromatic polycyclic hydrocarbons introduced during processing and have been reported to act as additional luminescent centers in some hBN samples Neumann et al. [2023]. These emitters contribute their own PL features and complicate the spectral response of the nanoparticles. Temporal instabilities such as blinking and bleaching, which are common in nanoscale fluorophores due to charge trapping or non-radiative recombination, further add to spectral variability. These intrinsic and extrinsic effects account for the observed unpredictability in the PL spectra of hBN nanoparticles and highlight the challenges of achieving reproducible optical performance.